OpenStudy (anonymous):

find the equation of the straight line which makes equal intercepts on the axes and passes through points (2,-7)

OpenStudy (anonymous):

@hartnn please help

hartnn (hartnn):

sure

OpenStudy (anonymous):

:)

hartnn (hartnn):

the equation of line can be y=mx+c do u know how to get the intercepts?

OpenStudy (anonymous):

yeah! If we need to find y intercept then we'll take x=0

hartnn (hartnn):

so find x and y intercepts for me.

OpenStudy (anonymous):

for that I'll have to find m first How would I do that?

hartnn (hartnn):

put x=0 in y=mx+c u get y=c<-----y-intercept. put y=0 in y=mx+c 0=mx+c mx=-c x= -c/m <----x-intercept

OpenStudy (anonymous):

and?

hartnn (hartnn):

but did u get that? now the line makes equal intercepts, so you have x intercept= y-intercept c= -c/m can u find m from here ?

OpenStudy (anonymous):

yep m= -1

hartnn (hartnn):

absolutely! (if you want you can remember that slope of line with equal intercepts is always -1, or you can do this exercise) so now your equation is y=-x+c ok?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

u have point (x,y)=(2,-7) on it so just substitute x=2 and y=-7 in y=-x+c and find c, try it

OpenStudy (anonymous):

I did that but what would be the equation?

hartnn (hartnn):

what is c ?

OpenStudy (anonymous):

-5

hartnn (hartnn):

-7=-2+c c= -5, correct. u had slope m=-1, u got c= -5 just put it in y=mx+c

OpenStudy (anonymous):

ohk and here too m= -1 right?

hartnn (hartnn):

its the same line! yes, slope = m= -1

OpenStudy (anonymous):

x + y + 5 =0

hartnn (hartnn):

that is correct! did u understand everything?

OpenStudy (anonymous):

yeah thanks a lot! You are an amazing tutor ;)

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

thanks for the compliment :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Refer to the attachment.

hartnn (hartnn):

how did u get that graph @robtobey ,the equation and intercepts are incorrect there.

hartnn (hartnn):

how is m=1 ?

hartnn (hartnn):

in your graph, u have y-intercept=-9 and x-intercept=9 so the intercepts are only equal in magnitude, if the question intends to have intercepts of equal magnitude, the there will be two lines as solution, one yours y=x-9, and one what we have derived, x+y+5=0. since the question asks for only one line, we assume that both intercepts must be equal in magnitude and sign, which gives the equation as x+y+5=0 only.

OpenStudy (anonymous):

Well the attachment below should do the trick, from the Mathematica program.